OpenStudy (anonymous):

24+-2 19/20+25 3/4 simplify, write as a proper fraction or whole number

OpenStudy (binary3i):

114/5

OpenStudy (diyadiya):

\[24+(-2) \frac{19}{20} =22 \frac{19}{20} = \frac{22 \times 20 +19}{20}= \frac{440+19}{20}= \frac{459}{20}\] \[25 \frac{3}{4}= \frac{25 \times 4 +3 }{4}= \frac{100+3}{4}= \frac{103}{4}\] \[\frac{459}{20}+ \frac{103}{4}= \frac{459}{20}+ \frac{515}{20}\] \[\frac{515+459}{20}= \frac{974}{20}= \frac{487}{10}= 48 \frac{7}{10}\]

OpenStudy (anonymous):

diyadiya - how do you get the fraction line in the equation function?

OpenStudy (diyadiya):

frac{numerator}{denominator} :)

OpenStudy (anonymous):

ok thanks

OpenStudy (diyadiya):

Welcome :)

OpenStudy (anonymous):

\[\frac{x+4}{3x-1}\]

OpenStudy (diyadiya):

:D

OpenStudy (anonymous):

ty!